OpenStudy (anonymous):

A singing competition is auditioning 47 contestants each hour for the next h hours. Which equation can be used to find the total number of contestants, C, that audition in h hours?

OpenStudy (anonymous):

A. h/47=C B.47+h=C C.47/h=C D.47h=C

OpenStudy (anonymous):

@skullpatrol